OpenStudy (anonymous):

Please help! Will award medal! lim x-->infinity of (9x^4 +3) / 4x^4 I could really use someone to guide me!

OpenStudy (anonymous):

\[\frac{ 9x ^{4}+3 }{ 4x ^{4} }=\frac{ 9 }{ 4 }+\frac{ 3 }{ 4x ^{4} }\]

OpenStudy (anonymous):

am I trying to get the denominators to be equal now?

OpenStudy (anonymous):

now substitute by x=infinity

OpenStudy (anonymous):

What does that mean? I am so sorry, my book is doing a crap job of explaining this

OpenStudy (anonymous):

now you have \[ \lim_{x \rightarrow \infty}(\frac{ 9 }{ 4 }+\frac{ 3 }{ 4x ^{4} })\]

OpenStudy (anonymous):

\[=\frac{ 9 }{ 4 }+\frac{ 3 }{ 4*\infty }\]

OpenStudy (anonymous):

how would I plug that in to a calculator?

OpenStudy (anonymous):

The limit is \[\infty\] isn't it?

OpenStudy (anonymous):

3/4*infinty =0